Solve by graphing
5x+2y=8
X-y=3

Answer by mananth(16949)

5 x + 2 y = 8
y= -5/ 2x +4
Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 4 2
y= 4 -6 -1
(x,y) (0,4),(4,-6),( 2-1)
--------
1 x + -1 y = 3
y= 1 x + -3

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 4 2
y= -3 1 -1
(x,y) (0,-3),(4,1),( 2,-1)


The solution is (2,-1)